can you please tell me how to find the square root of a number manualy without a calcalater. thanks
Answer by checkley71(8403)

I'LL TRY BUT IT WON'T BE EASY
---------3--2
---------------------------
-------V10 24 SQRT OF 1024 FIND THE LARGEST SQUARE NUMBER LESS THAN OR=10 (3)
---------9 3*3=9 SUBTRACT 9 FROM 10 WE GET 1
--------------
---------1 24 BRING DOWN THE NEXT 2 DIGITS & DOUBLE THE 3 (2*3=6)
-----62--1 24 DIVIDE 6 INTO 12 & WE GET 2 THEN MULTIPLY 32 BY 2 FOR 124
---------------
------------0 SUBTRACT 124 FROM 124 WE GET 0
32*32 OR 32~2=1024